### v1

# Facilitating Multi-Role and Multi-Behavior Collaboration of Large Language Models for Online Job Seeking and Recruiting

Abstract.
The emergence of online recruitment services has revolutionized the traditional landscape of job seeking and recruitment, necessitating the development of high-quality industrial applications to improve person-job fitting. Existing methods generally rely on modeling the latent semantics of resumes and job descriptions and learning a matching function between them. Inspired by the powerful role-playing capabilities of Large Language Models (LLMs), we propose to introduce a mock interview process between LLM-played interviewers and candidates. The mock interview conversations can provide additional evidence for candidate evaluation, thereby augmenting traditional person-job fitting based solely on resumes and job descriptions. However, characterizing these two roles in online recruitment still presents several challenges, such as developing the skills to raise interview questions, formulating appropriate answers, and evaluating two-sided fitness.
To this end, we propose MockLLM, a novel applicable framework that divides the person-job matching process into two modules: mock interview generation and two-sided evaluation in handshake protocol, jointly enhancing their performance through collaborative behaviors between interviewers and candidates. We design a role-playing framework as a multi-role and multi-behavior paradigm to enable a single LLM agent to effectively behave with multiple functions for both parties. Moreover, we propose reflection memory generation and dynamic prompt modification techniques to refine the behaviors of both sides, enabling continuous optimization of the augmented additional evidence. Extensive experimental results show that MockLLM can achieve the best performance on person-job matching accompanied by high mock interview quality, envisioning its emerging application in real online recruitment in the future.
Person-job fitting, Mock interview generation, Role-playing LLMs
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
With the emergence of online services, the job seeking and recruitment landscape has constantly been revolutionized in the job market. Across geographical boundaries, a large number of job seekers and recruiters now have greater opportunities to obtain suitable job positions and candidate talents through extensive contacts. For job seekers, they can apply for job interviews advertised in various job postings to seek positions that align with their skills and preferences. As to recruiters, they are responsible for assessing each candidate’s overall capabilities and making judgments based on job requirements. Online recruitment services have facilitated the acquisition of real-world interview data and further increased the need to develop higher-quality industrial applications for enhanced person-job matching.
Since candidate and recruiter information is usually represented as textual resumes and job descriptions, existing studies on person-job matching generally focus on modeling the latent semantics of resumes and job descriptions and learning a matching function between them to measure their suitability (Zhu et al., 2018a; Qin et al., 2018). Given the remarkable capabilities of Large Language Models (LLMs) in various tasks and their successful application in role-playing scenarios (Brown et al., 2020a; Ouyang et al., 2022a; OpenAI, 2023), it would be interesting to leverage LLMs to play both roles of interviewers and candidates through mock interviews, which are basically conversational interactions. These mock interview questions and answers can serve as augmented data to provide additional evidence for the evaluation process of both recruiters and job-seekers to facilitate better matching results.
However, existing role-playing frameworks typically assign each LLM role to specialize in a single task, optimizing its capabilities in such a specific aspect (Chen et al., 2023a, b). While for online recruitment, the interview conversation interactions and post-interview evaluation should be conducted by the same interviewer role to guarantee criteria consistency, and the same standard applies to the candidate role. It is natural for both roles to behave in multiple functions with collaborations. Moreover, the quality of mock interviews is critical but challenging. Conducting mock interviews requires high professionalism from interviewers by asking appropriate questions and making overall assessments. Similarly, it is also challenging for job seekers to provide decent answers and job hunting decisions. Hence, there are technical details to be solved during interviews and evaluations.
Figure 1. An example platform for interaction and evaluation between interviewers and candidates.
To this end, we propose MockLLM, a novel applicable framework to facilitate multi-role and multi-behavior collaboration for person-job matching. In contrast with the traditional role-playing framework in which the LLM agent is restricted to a single function, we design multiple behaviors for LLM-simulated interviewers and job seekers, enhancing the agent capabilities to handle various sub-tasks. The behaviors primarily undertaken by interviewers in the proposed method include: (1) raising interview questions in the interview stage: eliciting key information from the candidate’s resume and asking valuable questions via a multi-turn interview process; (2) evaluating interview performance in the review stage: assessing the appropriateness of candidates for the position based on their responses. On the other side, candidates engage with the following behaviors: (1) answering interview questions in the interview stage: providing qualified responses to the questions; (2) evaluating job positions: assessing the will of accepting an offer of the particular job position. To consider two-sided matching, we design a handshake protocol in MockLLM, integrating two-sided evaluation results to make final matching decisions. Figure 1 illustrates an example where interviewers and candidates follow the handshake protocol through two-sided evaluation for the matching in both directions.
Recruiters can interview multiple candidates, and vice versa, candidates will interview for multiple positions. To continuously enhance their behaviors and improve the quality of mock interviews, we propose reflection memory generation for both interviewers and candidates to memorize valuable interview experiences and optimize interviewing strategies. Once both parties reach a two-sided matching in the handshake protocol, the questions and answers in the interview are regarded as positive samples relevant to the candidate’s resume. For interviewers, successfully matched cases are memorized to face new candidates with similar backgrounds. For candidates, positive job positions are memorized to answer similar questions. The retrieved reflection memory in new interviews can be referred to optimize the corresponding prompts.
To sum up, our contributions can be summarized as manifold:
$\bullet$ We propose MockLLM, a novel applicable framework that leverages LLM role-playing capabilities to form mock interviews between interviewers and candidates, which can be utilized as augmented evidence for better evaluation of person-job fitting.
$\bullet$ We design the role-playing model within a multi-role and multi-behavior collaboration paradigm, enabling a single LLM role to master multiple behaviors of interviewers and candidates in mock interview generation and two-sided evaluation.
$\bullet$ We scheme reflection memory generation and dynamic prompt optimization techniques to refine the behaviors of interviewers and candidates. The strategies contribute to continuously improving the mock interview experiences to facilitate better person-job fitting.
2. Related Work
2.1. Person-Job Matching
Recruitment-oriented talent science research has consistently held a central position in human resource management, contributing to the prosperous growth of both enterprises and talents (Harris, 2017; Rendle et al., 2012). Person-job matching serves as the primary subject of recruitment-oriented talent science, which has witnessed continuous development and innovation in its research field (Zhu et al., 2018b).  (Malinowski et al., 2006) presents a recommendation model in the recruitment scenario through a personalized search for candidates. In the real job seeking and recruitment process, a two-way selection mechanism is essential for ensuring that both job seekers and recruiters find the best fit for their respective needs. NCF (He et al., 2017) is a collaborative filtering based method that utilizes multilayer perception to simulate the interaction between jobs and candidates. PJFNN (Zhu et al., 2018a) employs two convolutional neural networks to encode jobs and resumes separately, and then compute cosine similarity based matching scores. APJFNN (Qin et al., 2018) applies recurrent neural networks as document encoders and integrates attention mechanisms to model work abilities and skills. IPJF (Le et al., 2019) leverages multiple labels to indicate the propensity of candidates and jobs to reach a match.
In addition to job descriptions and resume texts, some research has found that it could bring benefits by incorporating interview history into person-job matching. (Yan et al., 2019) combines the interview history of job seekers and recruiters into the memory network to help improve person-job matching. JLMIA (Shen et al., 2018) leverages an unsupervised learning framework to integrate interview question generation and person-job fit using latent topic discovery. More recently, EZInterviewer (Li et al., 2023a) improves job interview quality using a knowledge-grounded dialogue model to connect the multi-turn interview session with candidates’ resumes. In this paper, we divide the person-job matching process into mock interview generation and two-sided evaluation in handshake protocol, jointly enhancing the performance of both modules to facilitate better person-job matching results.
2.2. Capabilities of Role-Played LLMs
With recent enhancements in model scales (Chowdhery et al., 2022; Ouyang et al., 2022b), LLMs have shown impressive capabilities in various domains such as text generation (Bubeck et al., 2023; Brown et al., 2020b), ranking (Ma et al., 2023; Chuang et al., 2023), and evaluation/verification (Shinn et al., 2023; Madaan et al., 2023). These powerful capabilities provide the opportunity to carefully design prompts to drive LLMs to play specific roles and complete specific tasks. CharecterLLM (Shao et al., 2023) gatherers character profiles from Wikipedia and teaches LLMs to act as specific people. (Wang et al., 2023) proposes RoleLLM, a framework to benchmark, elicit, and enhance role-playing abilities in LLMs. MetaGPT (Hong et al., 2023) is a multi-agent framework for assigning different roles to GPTs to form a collaborative software entity for complex tasks. It is a specialized LLM-based multi-agent framework for collaborative software development. CAMEL (Li et al., 2023b) is an LLM-based communicative agent framework that demonstrates how role-playing can be used to enable chat agents to communicate with each other for task completion.
However, traditional multi-role framework typically make a single LLM role/agent complete a specific sub-task, which is not in line with the requirements of interviewers and candidates. In this paper, we extend the existing LLM role-playing framework to a multi-role and multi-behavior collaboration paradigm. By assigning multiple behaviors for interviewers and candidates, they can generalize their capabilities in better simulation for job seeking and recruitment and promote better person-job matching.
3. MockLLM Framework
In this section, we will formulate our definitions and notations for the person-job matching process, and then introduce the proposed framework MockLLM in detail.
3.1. Problem Formulation
The interactions between job seekers (a.k.a., candidates) and job positions in the online recruitment data exhibit a two-sided one-to-many structure. Each candidate generally applies for multiple job positions, and each position, represented by a recruiter (a.k.a., interviewer), is responsible for assessing numerous candidates. Each candidate and job is accompanied by a descriptive text: a resume documents the candidate’s working skills and experiences, while a job description outlines job requirements and responsibilities of the position. In this paper, we divide the person-job matching process into two primary modules: (1) mock interview generation and (2) two-sided evaluation in handshake protocol.
Mock interview generation aims to simulate the multi-turn interview conversation between each interviewer $i_{k}$ and candidate $c_{l}$, where $k$ and $l$ are used as the indices of interviewers and candidates, respectively. Given an interview dialogue context $U_{n}=\{q_{1},a_{1},\cdots,q_{n},a_{n}\}$ containing $n$ historical question-answer pairs, the interviewer $i_{k}$ aims to pose a valuable and coherent question $q_{n+1}$ that aligns with the context $U_{n}$ and the candidate’s resume. In response, the candidate $c_{l}$ is expected to generate an insightful answer $a_{n+1}$ that demonstrates their competencies.
Once each interview is completed, a two-sided evaluation is carried out by both interviewer $i_{k}$ and candidate $c_{l}$ to determine the suitability of the candidate for the job position. Instead of traditional methods that solely rely on comparing the resume $r_{l}$ and job description $j_{k}$ to learn a matching function, we propose incorporating the mock interview dialogue history $U^{(i_{k},c_{l})}$ as a critical factor in person-job matching. Consequently, the final matching decision can be performed through a handshake protocol that considers the two-sided selection preferences, addressing the respective needs of both parties. Furthermore, we incorporate reflection to summarize the essential interviewing experience and use the reflection memory $M^{(i_{k})}$ and $M^{(c_{l})}$ to dynamically modify the interview behaviors for both parties, improving their ability to refer to previous successful experience when facing new similar interviews.
Figure 2. The framework overview of our MockLLM, which mainly consists of three modules: (1) Mock Interview Generation: The LLM role-played interviewer and candidate form mock interview through multi-turn conversations; (2) Two-Sided Evaluation in Handshake Protocol: Both parties evaluate each other based on the interview dialogue history, resume, and job description, and reach a matching consensus through a handshake protocol; (3) Reflection Memory Generation: Both roles save successfully matched cases to the reflection memory and apply them to modify questioning and answering prompts in subsequent interviews.
3.2. System Overview
In this section, we introduce the system overview of the proposed MockLLM, as illustrated in Figure 2. First, we customize the different roles in the online recruitment data. For each job description $j_{k}$, we instruct an LLM to comprehend the job requirements of $j_{k}$ and make it role-play an interviewer $i_{k}$ to communicate with job seekers and assess their performance. Similarly, we prompt an LLM to play a candidate $c_{l}$ based on the key information (e.g., skills and experiences) in each resume $r_{l}$ in response to the questions from the interviewer. In the following subsections, we will use $f_{*}(\cdot)$ and $g_{*}(\cdot)$ as the functions for interviewer $i_{k}$ and candidate $c_{l}$, respectively. Thus, the role-playing process can be given by:
(1) $i_{k}$ $=f_{role}(j_{k}),$
(2) $c_{l}$ $=g_{role}(r_{l}).$
To better formulate the person-job matching process, the proposed framework can be divided into two primary stages corresponding to the aforementioned modules: (1) the interview stage for mock interview generation, and (2) the post-interview review stage for two-sided evaluation in handshake protocol. These two stages are performed alternately, making the job seeking and recruitment scenario a continuous process. To enhance the ability of a single LLM role to perform multiple functions, we propose a multi-role and multi-behavior collaboration paradigm to enable interviewers and candidates to effectively complete tasks at each stage of person-job matching. For interviewers, their behaviors encompass interview question raising during the interview stage, and interview performance evaluation along with reflection memory generation during the review stage. As to job seekers, their behaviors include interview response generation in the interview stage, and job position evaluation as well as reflection memory generation in the review stage. Further details about these behaviors will be elaborated in subsequent subsections.
3.3. Mock Interview Generation
During the interview stage, interviewer $i_{k}$ and candidate $c_{l}$ engage in multiple turns of interaction in the conversation session. The behavioral norms of both parties in this stage are crucial, as the quality of the interviewer’s questions and the candidate’s answers directly affects the usability of the mock interview. They also serve as the basis for the evaluation of each other by both parties in the review stage. Therefore, to ensure the quality of the mock interview conversation, the goal of this module is to specify clear standards to guide the behaviors of interviewer $i_{k}$ and candidate $c_{l}$, controlling the generation of questions and responses. This can be divided into interview question raising for the interviewer and interview response generation for the candidate.
Interview question raising. Given an interview dialogue context $U_{n}=\{q_{1},a_{1},\cdots,q_{n},a_{n}\}$ comprising $n$ historical question-answer interaction pairs, interviewer $i_{k}$ is responsible for raising a new question $q_{n+1}$ in the $n+1$ turn. Due to the scarcity of expert resources, the LLM-played interviewers may not always prepare sufficiently valuable questions for systematically assessing candidates’ competencies, particularly in the context of person-job matching. To augment the professionalism of the interviewer, we establish well-defined standards for the question raising prompt $\theta_{q}$ concerning the posed question $q_{n+1}$: (1) coherence with the dialogue context $U_{n}$; (2) relevance to the job description $j_{k}$ and the candidate’s resume $r_{l}$; and (3) diversity among previous questions $\{q_{1},\cdots,q_{n}\}$ in $U_{n}$ to preclude repetition. In this way, we represent $f_{ques}$ as the question raising function, equipping all condition factors into $\theta_{q}$, and the newly raised question is thus formulated as:
(3) ${q}_{n+1}=f_{ques}(U_{n},r_{l},j_{k};\theta_{q}).$
Interview response generation. As the behaviors of both roles are tightly linked through multi-turn dialogue interactions, the primary objective of candidate $c_{l}$ in the interview stage is to generate an accurate answer $a_{n+1}$ in response to the question $q_{n+1}$. Concurrently, the candidates’ responses serve as a crucial factor in the subsequent two-sided evaluation and ultimate person-job matching. We also clarify the response generation standards in the answering prompt $\theta_{a}$, specifying that the provided answer $a_{n+1}$ in response to $q_{n+1}$ should be coherent with the entire updated dialogue context $\widetilde{U}_{n}=\{q_{1},a_{1},\cdots,q_{n},a_{n},q_{n+1}\}$. Additionally, the candidate’s insights and competencies described in the response must maintain relevance to the job description $j_{k}$ and the candidate’s resume $r_{l}$. The response generation process can be expressed by the function $g_{resp}$, as shown in the following equation:
(4) ${a}_{n+1}=g_{resp}(\widetilde{U}_{n},r_{l},j_{k};\theta_{a}).$
3.4. Two-Sided Evaluation in Handshake Protocol
The post-interview review stage immediately commences once an interview is concluded. The primary objective of this stage is to determine whether the candidate and the job are matched based on their respective evaluations of each other. Unlike traditional person-job fitting methods computing matching scores between resumes and job descriptions (Zhu et al., 2018a; Qin et al., 2018), the interview dialogue history provides alternative evidence for person-job matching decision-making. We believe that modeling the two-sided preference selection of interviewers and candidates yields the most accurate person-job matching results. Specifically, we integrate the interview dialogue history $U^{(i_{k},c_{l})}$ with the resume $r_{l}$ and job description $j_{k}$ to conduct a two-sided evaluation. This process can be divided into an interview performance evaluation for interviewer $i_{k}$ and a job position evaluation for candidate $c_{l}$. Then the proposed framework employs a handshake protocol based on the evaluation scores of both parties, forming a match only if both parties express mutual acceptance willingness.
Interview performance evaluation. From the perspective of interviewer $i_{k}$, to evaluate each candidate $c_{l}$ more accurately and systematically, it can be considered to divide the evaluation into two scoring steps. The interviewer first needs to give a basic fitting score $S_{r}^{(i_{k}\rightarrow c_{l})}$ for descriptive texts by comparing whether the skills and experience in the resume $r_{l}$ are suitable for the job requirements in the job description $j_{k}$. More importantly, the interviewer $i_{k}$ scores the candidate $c_{l}$’s performance $S_{q}^{(i_{k}\rightarrow c_{l})}$ in the interview dialogue history $U^{(i_{k},c_{l})}$ to determine whether $c_{l}$’s response matches the story in $c_{l}$ and meets the job requirements in $j_{k}$. Thus, the final evaluation score from $i_{k}$ to $c_{l}$ is related to a weighted average of $S_{r}^{(i_{k}\rightarrow c_{l})}$ and $S_{q}^{(i_{k}\rightarrow c_{l})}$, which can be formulated as:
(5) $S_{r}^{(i_{k}\rightarrow c_{l})}$ $=f_{eval}(r_{l},j_{k}),$
(6) $S_{q}^{(i_{k}\rightarrow c_{l})}$ $=f_{eval}\left(U^{(i_{k},c_{l})};r_{l},j_{k}\right),$
(7) $E^{(i_{k}\rightarrow c_{l})}$ $=\sigma\left(w_{r}\ S_{r}^{(i_{k}\rightarrow c_{l})}+w_{q}\ S_{q}^{(i_{k}\rightarrow c_{l})}\right),$
where $\sigma(\cdot)$ denotes the sigmoid function.
Job position evaluation. The evaluation behavior from the candidate’s perspective is dual to that of the interviewer. First, candidate $c_{l}$ needs to assign a basic fitting score $S_{j}^{(c_{l}\rightarrow i_{k})}$ between his/her own resume and the requirements in job description $j_{k}$. Moreover, combining the interview dialogue history $U^{(i_{k},c_{l})}$, the candidate $c_{l}$ should further evaluate whether the position is suitable for him and score the opportunity of obtaining the offer for the position. The final evaluation from $c_{l}$ to $i_{k}$ can also be viewed as a weighted average of $S_{j}^{(c_{l}\rightarrow i_{k})}$ and $S_{a}^{(c_{l}\rightarrow i_{k})}$, which is given by:
(8) $S_{j}^{(c_{l}\rightarrow i_{k})}$ $=g_{eval}(r_{l},j_{k}),$
(9) $S_{a}^{(c_{l}\rightarrow i_{k})}$ $=g_{eval}\left(U^{(i_{k},c_{l})};r_{l},j_{k}\right),$
(10) $E^{(c_{l}\rightarrow i_{k})}$ $=\sigma\left(w_{j}\ S_{j}^{(c_{l}\rightarrow i_{k})}+w_{a}\ S_{a}^{(c_{l}\rightarrow i_{k})}\right).$
Person-job matching in hardshake protocol. Based on the comprehensive evaluation ensuring a more accurate and effective two-sided preference selection, the eventual decision of person-job making is performed by following a handshake protocol. Both the interviewer and the candidate have the right to choose whether to accept or reject each other, helping both parties better understand each other’s needs and abilities, and promoting better person-job matching. This handshake protocol can be formulated as:
(11) $y(i_{k},c_{l})=\left\{\begin{aligned} 1,&\quad\text{if }E^{(i_{k}\rightarrow c_{l})}=1\text{ and }E^{(c_{l}\rightarrow i_{k})}=1,\\ 0,&\quad\text{otherwise}.\end{aligned}\right.$
3.5. Reflection Memory Generation
The interaction during the interview stage provides a forward reference for the evaluation during the review stage. In order to better enable our framework to simulate real job seeking and recruitment scenarios, we consider feedback during the review stage to enhance the quality of the subsequent interview stage. Specifically, in addition to the two-sided evaluation behaviors of job matching, we design reflection behaviors for both parties during the review stage to summarize the experience in the interview and provide feedback to improve their behavior in subsequent interviews. For both interviewers and candidates, we first initialize the natural language based questioning and answering strategies, which are denoted as $m_{0}^{i}$ and $m_{0}^{c}$, respectively. Then, we consider leveraging the positive matching cases to continuously refine the strategies. Taking interviewer $i_{k}$ as an example, if candidate $c_{l}$ achieves a two-sided matching with him/her, then their interview dialogue history $U^{(i_{k},c_{l})}$ is considered to be of relatively high quality and has guiding value for recruiting new candidates in the future interviews. Then, we use a reflection function $f_{ref}$ to refine the questioning strategy to $m^{(i_{k}\rightarrow c_{l})}$, which can be more targeted to figure out the specific skills and experiences in the raised question. Then, we incorporate $c_{l}$ into the $i_{k}$’s reflection memory $M^{(i_{k})}$, including the tuple format: the resume $r_{l}$, the interview dialogue history $U^{(i_{k},c_{l})}$, and the refined questioning strategy $m^{(i_{k}\rightarrow c_{l})}$. Similar to each candidate $c_{l}$, the answering strategy can be refined to $m^{(c_{l}\rightarrow i_{k})}$ through $g_{ref}$.
(12) $m^{(i_{k}\rightarrow c_{l})}=\left\{\begin{aligned} &f_{ref}\left(r_{l},U^{(i_{k},c_{l})}\right),&\quad\text{if }y(i_{k},c_{l})=1.\\ &m_{0}^{i},&\quad\text{otherwise}.\end{aligned}\right.$
(13) $M^{(i_{k})}\leftarrow M^{(i_{k})}\bigcup\big{\{}c_{l}|r_{l},U^{(i_{k},c_{l})},m^{(i_{k}\rightarrow c_{l})}\big{\}}.$
(14) $m^{(c_{l}\rightarrow i_{k})}=\left\{\begin{aligned} &g_{ref}\left(j_{k},U^{(i_{k},c_{l})}\right),&\quad\text{if }y(i_{k},c_{l})=1.\\ &m_{0}^{c},&\quad\text{otherwise}.\end{aligned}\right.$
(15) $M^{(c_{l})}\leftarrow M^{(c_{l})}\bigcup\big{\{}i_{k}|j_{k},U^{(i_{k},c_{l})},m^{(c_{l}\rightarrow i_{k})}\big{\}}.$
3.6. Prompt Modification with Feedback
In order to further link the two stages of interview and review, we consider using the feedback from the review stage to help improve the quality of subsequent interviews. Specifically, the feedback is the memory formed by reflection. The specific optimization method is to modify the prompts for interview questions and responses to make them more targeted. The initial questioning prompt applies to all candidates, but after modification, the interviewer will ask questions more specifically based on the skills and experience of different interviewers. The same goes for job seekers. The prompts they reply to are initially short and general, but as the number of interviews increases, the interview experience is accumulated through the reflection process. The prompt optimization of questioning lies in two aspects: (1) Based on the summary content formed by reflection, the strategy for questioning is improved, so that the questions asked in different candidates and different rounds are more targeted; (2) Extracting similar interviews from memory Information allows the interviewer to find job seekers with similar job skills or experience that have been interviewed before in the memory, and use the questions they asked as a reference when asking new candidates.
(16) ${\theta}_{q}$ $\leftarrow f_{mod}(\theta_{q},M^{(i_{k})}_{c^{\prime}},r^{\prime}),$
(17) ${\theta}_{a}$ $\leftarrow f_{mod}(\theta_{a},M^{(c_{l})}_{i^{\prime}},j^{\prime}),$
4. Experimental Setup
4.1. Data
In this paper, we conduct experiments on the online recruitment datasets sourced from “Boss Zhipin” [^1] https://www.zhipin.com, the largest online recruiting platform in China. To protect the privacy of candidates, all textual documents are anonymized by removing all personal identifying information. The preprocessed dataset contains 1,992 resumes (candidates) and 1,968 job descriptions (interviewers) in total. More details of dataset statistics are shown in Table 1.
4.2. Evaluation Metrics
Following the previous studies (Zhao et al., 2022b), we employ four widely used top-$k$ recommendation metrics to test the effectiveness of person-job matching in each side of candidates or jobs. These metrics include Recall (R@k), Precision (P@k), Normalized Discounted Cumulative Gain (NDCG@k), and Mean Reciprocal Rank (MRR@k). To further evaluate two-sided matching performance, we use Macro Precision, Recall, F1 scores to calculate the accuracy of successful matching between both parties.
To evaluate the quality of mock interview generation, we use BLEU (B@n) for n-gram overlap between generated utterances and candidates’ resumes. In addition to these automatic metrics, we further perform a human evaluation for a more comprehensive assessment of the mock interview generation quality, incorporating the following metrics: (1) Coherence: The extent to which the generated questions/answers are coherent with the dialogue context; (2) Relevance: The degree to which the generated questions/answers are relevant to the candidate’s resume; (3) Diversity: The level of variation in the generated questions/answers compared to previous utterances in the dialogue history.
4.3. Comparisons
To evaluate the effectiveness of the proposed framework for person-job matching, we select the following methods for comparison, including (1) NCF (He et al., 2017) leverages a collaborative filtering method to model the interaction between jobs and candidates; (2) PJFNN (Zhu et al., 2018a) employs convolutional neural networks to encode resumes and job descriptions separately and computes the matching score using cosine similarity; (3) APJFNN (Qin et al., 2018) leverages semantic representations for both resumes and job descriptions based on sequential networks along with attention hierarchically; (4) IPJF (Le et al., 2019) uses the multi-level interactions as supervision signals to indicate the preferences for candidates and jobs.
Moreover, to assess the quality of mock interview generation, we choose classical specialized pre-trained models, such as GPT-2 (Radford et al., 2019) and T5-base (Raffel et al., 2020), in the TecentPretrain (Zhao et al., 2022a) toolkit as the backbone, and fine-tune them in our training set. Inspired by the knowledge-grounded dialogue generation methods, we use K2R (Adolphs et al., 2021) as a representative to first generate a knowledge sentence based on the context, and then produce a final response.
Table 1. Statistics of the datasets used in the experiments.
Statistics Values
# of resumes 1,992
# of job descriptions 1,968
# of interview conversations 61,215
Avg. turns per session 3.19
Avg. tokens per dialog 560
Avg. tokens per resume 1,553
Avg. tokens per job description 405
# of context-response pairs 195,318
4.4. Implementation Details
In the proposed framework, we take advantage of the role-playing capabilities of LLMs, forming augmented mock interviews to enhance the evaluation process. We employ the open-sourced general LLM Nanbeige-16b-chat (Lab, 2023) as our backbone to play the roles of interviewers and candidates. We use the vLLM (Kwon et al., 2023) as the inference framework to make LLMs efficiently generate outputs. We set the hyper-parameters in LLMs as follows: temperature=0.9, top_p=0.6, repetition_penalty=1.1, and max_tokens=32,768. We use the training set to update and accumulate reflection memory, and then directly perform inference on the test set without any additional fine-tuning. [^2] Our code is available at https://github.com/p1nksnow/PJF_interview. For the implementation of baseline models, we uniformly use the default fine-tuning hyper-parameters for TecentPretrain (Zhao et al., 2022a) and K2R (Adolphs et al., 2021) and set the learning rate as 0.001.
5. Experimental Results
5.1. Overall Performance
We first discuss the performance comparison results of two-sided person-job matching. As presented in Table 2, we can see that content-based methods, such as APJFNN and PJFNN, perform less well than collaborative filtering-based NCF. This could be because the matching approaches in APJFNN and PJFNN are case-specific and highly dependent on the representation of textual documents. They potentially struggle to fully comprehend the organizational structure of resumes and job descriptions in our dataset. In contrast, NCF, which can leverage similar matching cases to infer the compatibility of a new case, exhibits slightly better performance. Moreover, the hybrid-based method IPJF uses both text representation and information from similar cases to yield improved matching results. Our proposed MockLLM achieves the best matching accuracy for jobs, candidates, and both sides. We incorporate mock interview conversations to help enrich the evaluation stage. Meanwhile, the handshake protocol in the proposed framework is suitable for two-sided matching, which supports being put into practical applications to allow interviewers to effectively screen suitable candidates and enable candidates to find appropriate positions.
Table 2. Performance comparison of two-sided person-job matching. Bold numbers indicate that we accept the hypothesis of model improvement relative to the best baseline at a significance test level of 0.01.
Direction Candidates Jobs Both Sides
Method NDCG@5 R@5 P@5 MRR@5 NDCG@5 R@5 P@5 MRR@5 Precision Recall F1
NCF (He et al., 2017) 0.296 0.356 0.207 0.401 0.273 0.329 0.193 0.375 0.508 0.516 0.425
APJFNN (Qin et al., 2018) 0.269 0.324 0.186 0.378 0.258 0.314 0.188 0.348 0.501 0.502 0.431
PJFNN (Zhu et al., 2018a) 0.275 0.332 0.192 0.374 0.260 0.318 0.187 0.356 0.503 0.505 0.453
IPJF (Le et al., 2019) 0.353 0.427 0.252 0.440 0.340 0.420 0.250 0.441 0.549 0.596 0.477
MockLLM w/o MI 0.218 0.297 0.167 0.236 0.199 0.258 0.146 0.223 0.591 0.622 0.578
MockLLM w/o SU 0.366 0.450 0.261 0.422 0.376 0.463 0.271 0.432 0.597 0.659 0.583
MockLLM w/o RP 0.372 0.453 0.262 0.430 0.380 0.470 0.277 0.434 0.590 0.651 0.560
MockLLM 0.395 0.479 0.279 0.443 0.413 0.494 0.296 0.475 0.611 0.685 0.586
Table 3. Automatic and human evaluation results of mock interview generation.
Method Questions Answers Human Eval
B@1 B@2 B@3 B@4 B@1 B@2 B@3 B@4 Coherence Relevance Diversity
TencentPretrain (GPT2) 13.74 8.76 5.45 3.40 2.53 1.23 0.58 0.27 2.26 2.05 2.08
TencentPretrain (T5-base) 17.70 13.53 10.58 8.52 2.52 1.48 0.89 0.59 3.21 3.34 3.15
TencentPretrain+K2R (GPT2) 14.24 8.50 4.74 2.62 2.72 1.33 0.60 0.27 2.34 1.99 2.16
TencentPretrain+K2R (T5-base) 18.36 13.35 9.92 7.68 2.96 1.70 0.99 0.62 3.45 3.37 3.23
MockLLM 24.23 16.75 12.14 9.20 16.48 9.87 6.36 4.53 4.12 4.03 4.20
5.2. Ablation Study
To verify the effectiveness and necessity of various components of the proposed framework, we conduct a comparison using the following ablation variants: (1) MockLLM w/o MI eliminates the connections between mock interview generation and evaluation, making the evaluation process rely solely on resumes and job descriptions; (2) MockLLM w/o RP removes the reflection behaviors and prompt modification, and only retains the two-stage forward process of person-job matching; (3) MockLLM w/o SU omits the strategy updates and consistently uses the initial strategies $m_{0}^{i}$ and $m_{0}^{c}$ for all mock interviews.
The ablation results are also shown in Table 2. The performance of MockLLM w/o MI degrades most significantly, indicating the necessity of the proposed framework that integrates the mock interview dialogue to support talent and job evaluation and the final matching decision-making process. The performance gap between MockLLM w/o RP and the full model demonstrates the effectiveness of the reflection memory generation module in facilitating the continuous refinement of interviewer and candidate behavior. MockLLM w/o SU is unable to generate targeted and valuable questions or answers, which confirms the importance of strategy updates to the framework.
5.3. Impact of Mock Interview Generation
As an innovative module of our framework, the quality of mock interview generation directly influences the matching evaluation criteria, so it also needs to be evaluated.
Automatic evaluation. As shown in the left part of Table  3, we can observe that the specially fine-tuned dialogue model is still not as effective as MockLLM in mock interview generation. After K2R uses generated outcomes as an intermediate reference to generate final utterances, the performance of the same backbone can be slightly improved in relevance to resumes and job descriptions. Our MockLLM can leverage the LLM role-playing capability to cooperate with interviewers and candidates. The multi-role and multi-behavior collaboration framework can indeed improve the quality of mock interview generation, making interviews more centered around resume content and job requirements.
Human evaluation. We employ three well-educated volunteer annotators to further evaluate the quality of questions and responses generated by the mock interviews. The evaluation is performed in a double-blind manner ensuring that the annotators do not know which output corresponds to which method. We randomly sample 100 cases from the test set and ask each annotator to rate the questions and responses generated by each method in the aforementioned metrics: Coherence, Relevance, and Diversity.
The comparison results of human evaluation are listed in the right side of Table 3. Compared to baseline methods, our proposed MockLLM can generate more coherent questions and responses to align with the interview context. The improvement on the Relevance score indicates that the questions raised by MockLLM are more relevant to the candidate’s resume, making the topics discussed in the interview conversation accurate and valuable. In addition, the highest Diversity means that MockLLM can raise more questions from new directions based on existing questions, reducing the repetition of questions to ensure the quality of mock interviews.
Figure 3. An example of mock interview generation and two-sided evaluation shows the superiority of MockLLM, which can generate higher-quality mock interviews and achieve better two-sided person-job matching decisions.
5.4. Impact of Number of Retrieved Cases
During the prompt modification process, we need to retrieve relevant cases from reflection memory and equip them as references into questioning/answering prompts. Therefore, we conduct this analysis of the suitable number of retrieved cases in our framework. We prompt LLMs to retrieve the top-$k$ relevant cases ($k=0,1,2,3$) from reflection memory. The results in Figure 5 show that incorporating the historical experience summarized in the reflection memory is truly helpful to continuously enhance the abilities of interviewers and candidates. However, retrieving more cases does not always lead to better results. In fact, including too many cases may introduce noise that affects the interview quality.
Figure 4. Comparison results before and after questioning prompt modification.
5.5. Qualitative Analysis
In addition to the above quantitative analysis, we will provide sufficient cases to illustrate the specific applications of the proposed framework in the qualitative analysis in this subsection.
Case study. We present a person-job matching example for comparison with our proposed MockLLM and baseline models in Figure 3. From the given resume, we observe that the candidate is proficient in Java and Linux, and has experience in database management projects. The interview questions simulated by the interviewer in MockLLM not only match the candidate’s background but also enable targeted conversations revolved around their key skills and experiences. In contrast, baseline models fail to raise valuable questions that are sufficiently connected to the resume in multi-turn conversations. As to the evaluation process, only our method reaches the correct matching results through two-sided evaluation in handshake protocol, where baseline methods fail to make an accurate matching decision as they cannot incorporate mock interview history into the evaluation process.
Analysis of prompt modification. Regarding our proposed prompt modification technique, we provide specific results before and after questioning prompt modification and give more elaborations for qualitative comparative analysis. Figure 4 shows that the initial questioning prompt without modification is a general instruction for all candidates. Compared to the initial strategy, the modified prompt expands more questioning requirements and insights in detail and incorporates relevant cases from reflection memory to draw on prior experience to help raise new valuable questions.
Figure 5. The impact of the number of retrieved cases.
6. Conclusion
In this paper, we propose MockLLM, a novel industrial framework that leverages LLM role-play capabilities to form mock interview conversations between interviewers and candidates, which serves as augmented evidence for better evaluation of person-job matching. We extend the role-playing scenario to a multi-role and multi-behavior collaboration paradigm, enabling a single LLM role to master multiple behaviors of interviewers and candidates in mock interview generation and two-sided evaluation. We utilize reflection memory generation and dynamic prompt optimization techniques to refine the behaviors of interviewers and candidates. This contributes to continuously improving the interviewer’s professionalism and the candidate’s interview experience to facilitate better person-job matching.